OpenStudy (anonymous):

The resistance of a thermistor, a device used to monitor the temperature inside an engine, changes with temperature. The resistance of the thermistor is given by R(t) = t (third power) + 5, where R is measured in ohms and t is the temperature in oC. What is the average rate of change of resistance per degree of temperature change between 0oC and 2oC? this is soo hard

OpenStudy (anonymous):

R(t) = t³ + 5 R'(t) = 3t² //Rate of change R'(0) = 0 R'(2) = 12 I'd say it is (12+0)/2 = 6

OpenStudy (anonymous):

ok here are the answer choices 4 8.5 11 17

OpenStudy (anonymous):

oh sh.t hahahaha

OpenStudy (anonymous):

lol what ?

OpenStudy (anonymous):

it's 4

OpenStudy (anonymous):

rate = f(2) - f(0) / 2 - 0 rate = 13 - 5 / 2 rate = 8/2 = 4